

********************************************************************
UARS WEEKLY STATUS REPORT - 05 FEB 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
data without apparent problems.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
PEM is active PMON control. Energy processor #6 for the HEPS-1 unit
has been found to not activate correctly. Data affected are from the upper
energy EE channel for Head-2 of HEPS-1 (that is the +15 deg electrons above
300 keV). Tests are being conducted on this energy processor to assure
that it is operating correctly. Measurements from this sensor will be
activated as soon as possible. The PEM team and FOT are currently examing
the PMON activation sequence to determine if there is interference during
the commanding of this energy processor.
While under PMON control, the PEM instrument computes energy deposition
from the HEPS unit.
***********
HALOE
***********
HALOE is scheduled to be turned on February 8, 1999. HALOE has been off
during this high beta period beginning January 25, 1999.
V19 is now the archived version of the HALOE data.
***********
MLS
***********
MLS was reactivated successfully on Jan 26 and is scheduled to operate through
Feb 11. The MLS team is operating the instrument in a mode emphasizing
scientific observations at the northern latitudes.
***********
HRDI
***********
HRDI continues to operate normally.
***********
WINDII
***********
WINDII is on and operating normally.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
106 days SUS (7.00) PROD (L3) (2051R-1845C-10F=196L)
0 days WIN (5.11) ENG (L1-L3) (1R-1F=1L)
1 day OBC (3.4) PROD (TEST) (1R-1C=0L)
1 day ATT (7.2) PROD (TEST) (1R-1C=0L)
The CDHF plans to run a SOLSTICE day that used excessive Wallclock time
with an Engineering Executable supplied by UCSS developer.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 January 1999 (Orbit 40256, GMT Day 023) through 29
January 1999 (Orbit 40361, GMT Day 029).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 023/0616 40260 PEM Test
HALOE 024/0937 40277 Off
MLS 026/0039 40302 On
WINDII 026/1319 40310 On
The beta angle ranged between 43.3 degrees to 69.7 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 74.0
degrees on 31 January 1999 (DOY 011).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 22 January 1999 and 29 January 1999 are provided in
Appendix D.
The clock error ranged from -3.0 to 20.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HALOE 01
% HRDI 02
% MLS 01
% PEM 04
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/362 - 2287.496912
Transponder B: 99/357 - 2287.499093
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/028 show the observatory data loss to be
21 hours, 50 minutes, 50.292 seconds (an increase of 29.504 seconds since
last report period). This is a 0.0339 percent data loss which equals a
99.9661 percent data capture for the mission.
The increase of 29.504 seconds occurred over two (2) days:
7 9.152 seconds on Day 023 due to insufficient overlap, and
7 20.352 seconds on Day 016 due to insufficient overlap, data hit (IP),
and data hit.
Reproduction of the FOT Operational Database #069 has been completed. Two
copies are resident Ron-lineS and the remainder are used by the Roff-lineS
engineers.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. One test of a modified command sequence failed
to start EP6. SWRI requested a new test sequence which will be loaded
early in FW 06.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Testing of the new health and safety system showed great strides since the
last release. Unfortunately, checkout tests of a candidate release
revealed that some nagging telemetry decoding errors remain. The FOT is
working with the sustaining engineering team to test and validate all
telemetry points on the Health & Safety (H&S) database pages.
MEETINGS
A UARS FOT meeting was held on 22 January 1999 to discuss the concepts or
theories for possible preventative actions to minimize the full Sun charge
impact on batteries in Sun Pointing Safehold mode.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 20
February 1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
Carlos Hester, UARS Ground Controller, transferred to the EOS-AM project
effective 29 January 1999.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 January 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40260 023/0616 99-033 PEM PEM TEST
40263 023/1115 99-034 PEM PEM TEST
40274 024/0431 99-035 OBC Clk Adj fr 0.00 to 20.6 msec/day
40277 024/0937 99-036 HALOE Power OFF
40295 025/1423 99-037 PEM RTS-10 Activate
40302 026/0039 99-038 MLS Power ON
40310 026/1319 99-039 WINDII Power ON
40326 027/1455 99-040 OBC Dump Table-74
40326 027/1609 99-041 PEM PDB-9 Up-date
40339 028/1235 99-042 OBC Clk Adj 20.6 to 0.00 msec/day
40431 028/1616 99-043 HRDI HRDI to Safehold
40343 028/1931 99-044 OBC Clk Adj 0.00 to 41.2 msec/day
40345 028/2209 99-045 HRDI HRDI Microloads
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 January 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HATBAT Y-HI HALOE TBAT TEMP 40276
WIQBCURRENT Y-LO WINDII Quiet Bus Current 40311-17,20-31,35-61
HRPTSBARTMP Y-LO HRDI BARREL TEMP 32.8 40336-39,43
HRPTSBARTMP R-LO HRDI BARREL TEMP 48.6 40341,42
QTZEPSIF1 Y-LO ZEPS I/F TEMP 40345
QTZEPSIF2 Y-LO ZEPS I/F TEMP 40345,47-8
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 January 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40274 024/04:34 (24 Jan 99) from 00.0 to 20.6 msec/day
40339 028/12:40 (28 Jan 99) from 20.6 to 00.0 msec/day
40343 028/19:42 (28 Jan 99) from 00.0 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 January 1999
22 Jan 1998 - GMT Day 022 Beta = 43.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +39.2/+11.2 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-50.4 +6.44/+4.51 +29.6/-16.0 25.1 1.023 22.1
3 +33.6/-11.2 +5.15/+3.87 +28.8/-16.0 25.1 1.026 21.8
29 Jan 1999 - GMT Day 029 Beta = 69.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +22.4/ 0.0 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-16.8 +3.87/+2.60 +20.8/-16.0 26.6 1.023 18.1
3 +39.2/-11.2 +2.92/+1.34 +20.0/-15.2 26.6 1.027 17.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 January 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-005 FOT 40310 026/1458 20535 Loss of signal. 13m 49sec
R/T data lost/recoverable
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-005 FOT 40310 026/1458 20535 Loss of signal. 13m 49sec
R/T data lost/recoverable
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-001 FOT see list Jan 98 14170 Generic Late Acq
SEVNTFW99/05